In an electrical circuit, the current passing through a conductor varies inversely with the resistance. Suppose that when the current is 10 A (amperes), the resistance is 14 ohms. What is the resistance when the current is 28 A?
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Inverse variation means the product of the current and resistance is constant. So

10*14 = 28*x
140 = 28x
x = 140/28 = 5

ANSWER: 5 ohms

-----------------------------------------------------------------

I strongly disagree with the response from the other tutor.

Since the problem states that the current varies inversely with the resistance, it is clear that the problem is NOT talking about the same circuit in both cases; the conductors in the two circuits are NOT the same.

All other factors the same, when the current in the two circuits are different, the resistances are different. When the current is 28A, the resistance is 5 ohms.


Answer by ikleyn(52809)

.

The other tutor doesn't know what a resistance is.

Under normal conditions, the resistance is the interior property of the conductor and DOES NOT DEPEND on the current or voltages.

Under given conditions, the resistance is 14 ohms in both cases, if the conductor is the same.

================

The wording of the problem is simply wild in its illiteracy, so I better will not comment it.